### v1

# On almost commuting unitary matrices

Abstract.
A question going back to Halmos asks when two approximately commuting matrices of a certain kind are close to exactly commuting matrices of the same kind. It has long been known that there is a winding number obstruction for approximately commuting unitary matrices to be close, in a dimension-independent way, to genuinely commuting unitary matrices. In this paper, under the vanishing of the said obstruction, we obtain effective bounds for the distance to commuting unitary matrices in terms of the commutator of the original matrices.
A. D. was partially supported by NSF / BSF grants 2350543 / 2023695 (respectively), 2452324 / 2024734 (respectively) and a DFG Middle-Eastern collaboration project 529300231.
L. H. was supported as a Zuckerman Postdoctoral Scholar in the 2023-2024 cohort.
I. K. was partially supported by the NSF grants DMS–1846114, DMS–2052519, and the 2022 Sloan Research Fellowship.
1. Introduction
Let $H$ be a (complex) Hilbert space. In 1976, P. Halmos stated the following question in [16]:
Question 1.1.
Is it true that for every $\varepsilon>0$ there is a $\delta>0$ such that, if $x,y\in B(H)$ satisfy
(1.1) $\|x\|\leq 1\quad\|y\|\leq 1\quad\|[x,y]\|=\|xy-yx\|<\delta,$
then there are commuting operators $x^{\prime},y^{\prime}\in B(H)$ satisfying
(1.2) $\|x-x^{\prime}\|+\|y-y^{\prime}\|<\varepsilon?$
It is common to refer to equation (1.1) as $x$ and $y$ being almost commuting, and equation (1.2) as being nearly commuting (cf. [8]), with the question itself stated as are almost commuting matrices nearly commuting?
As already pointed out in [16], this question has a negative answer in infinite dimensions [4]. However, one can produce many non-trivial versions of it by imposing various natural restrictions on $x$ and $y$, even if $\dim H<+\infty$. In particular, a large body of work focuses on the variants of Question 1.1 with the following conventions:
  • $\dim H<+\infty$, so that $x$ and $y$ can be considered to be finite matrices. However, $\delta$ is allowed to depend on $\varepsilon$ but not on the dimension of $H$ (“dimension-independent results”).
  • $\|\cdot\|$ is the usual operator norm.
  • The matrices $x$ and $y$ belong to a certain class. For example, as originally suggested in [16], one can consider Hermitian matrices $x=x^{*}$, $y=y^{*}$.
By considering $a=x+iy$, one can restate the self-adjoint version of the above question as whether the distance from $a$ to normal operators can be estimated in terms of $\|[a,a^{*}]\|$ (that is, the norm of the self-commutator of $a$). In this form, positive answer to the question was provided in [22]. See also [13] for a shorter proof. In both cases, it is reduced by contradiction to a lifting problem in a certain ultraproduct of matrix algebras, and extracting an explicit dependence relation between $\varepsilon$ and $\delta$ from this kind of proof becomes a challenging reverse engineering problem with unclear possibilities of obtaining a optimal bound. An alternative approach to the self-adjoint problem was proposed in [21], which led to an order sharp estimate $\operatorname{dist}(a,\textrm{\{normal matrices\}})\leq C\|[a,a^{*}]\|^{1/2}$, corresponding to $\delta=C\varepsilon^{2}$. See also [17, 19] regarding earlier results and [8] for a survey.
1.1. Almost commuting unitary matrices: overview and the main result
We will be mainly interested in Question 1.1 for a pair of unitary matrices (which will be denoted by $u$ and $v$ rather than $x$ and $y$). In this case, the problem presents a topological obstruction, see [25, 12, 11], which has been characterized in several somewhat equivalent ways, the most simple of them being the winding number invariant: for $u,v\in U(n)$ with $\|[u,v]\|<2$, the winding number $w(u,v)$ of the curve
(1.3) $\omega:[0,1]\rightarrow\mathbb{C}\setminus\{0\},\quad\omega(t):=\det(t\cdot uv+(1-t)\cdot vu).$
is well-defined. Assuming that $\varepsilon$ is small enough, it is shown in the above references that any pair $u,v\in U(n)$ that is close to a commuting pair must have $w(u,v)=0$. This is confirmed by the example of Voiculescu’s unitaries [25], a family of pairs of matrices $u_{n},v_{n}\in U(n)$ with $\|[u_{n},v_{n}]\|=O(1/n)$ whose distance to any commuting pair is bounded from below by $\sqrt{2}-1-o(1)$ as $n\to+\infty$ [11].
In view of the discussion above, a natural setting for Question 1.1 for two unitaries includes the additional assumption that their winding number invariant vanishes. In this form, [10] and [15] independently resolved the problem around the same time. Similarly to the self-adjoint case considered in [22, 13], both solutions involve the ultraproduct construction and therefore do not provide an explicit relation between $\varepsilon$} and $\delta$.
The present paper answers Question 1.1 for unitary matrices in the above setting with the following quantitative bounds:
Theorem 1.2.
There exists an absolute constant $C>0$ such that for any $u,v\in U(n)$ with $w(u,v)=0$ one can find $u^{\prime},v^{\prime}\in U(n)$ such that
$u^{\prime},v^{\prime}]=0,\quad\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq C\|[u,v]\|^{1/30}.$
1.2. Overview of the proof
First, we note that there are three versions of the winding number invariant, with different potential directions of generalizations; however, all three coincide in our setting of a pair of almost commuting unitary matrices. We will use two of them: the winding number $w(u,v)$ of the curve defined by equation (1.3) and the isospectral invariant $\operatorname{isospec}(u,v)$, originally introduced in [5, Theorem 4.1], see equation (3.2) in Section 3. For the convenience of the reader, we show that $w(u,v)=\operatorname{isospec}(u,v)$ in Corollary 3.4, since the original proof is spread among several of the above-referenced papers.
The proof of Theorem 1.2 is a combination of the following two results obtained in this paper. The first one provides quantitative estimates for the isospectral homotopy lemma from [5].
Lemma 1.3.
There exists an absolute constant $C>0$ such that for every two unitary matrices $u,v\in U(n)$ with $\operatorname{isospec}(u,v)=0$, there exists a piecewise smooth path $v_{t}\colon[0,1]\to U(n)$ of fixed length with
(1.4) $v_{0}=v,\quad v_{1}=\mathbf{1},\quad\|[v_{t},u]\|\leq C\|[u,v]\|^{2/5}.$
The second result gives a quantitative variant of a theorem, whose analogue in the setting of real rank zero $C^{*}$-algebras was proved in [23] and [24] independently and around the same time.
Theorem 1.4.
Suppose that $u,v$ are unitary matrices such that $\|[u,v]\|\leq\delta$ and there exists a continuous path $\{u_{t}\colon t\in[0,1]\}$ such that:
$u_{0}=u,\quad u_{1}=\mathbf{1},\quad\|[u_{t},v]\|\leq\delta\quad\forall t\in[0,1],$
Then there exist unitary matrices $u^{\prime}$, $v^{\prime}$ such that
$u^{\prime},v^{\prime}]=0,\quad\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq C\delta^{1/12}$
Proof of Theorem 1.2. As mentioned earlier, Corollary 3.4 implies that $\operatorname{isospec}(u,v)=0$. Then, the result follows from Theorem 1.4 applied with $\delta=C\|[u,v]\|^{2/5}$, using the conclusion of Lemma 1.3.  ∎
In the cases of Lemma 1.3 and Theorem 1.4, we focus on the setting of two unitary matrices. It is likely that Lemma 1.3 and 1.4 can be proved in the level of generality close to the original results [5, 23]. We intend to explore it in the future, as well as potential generalizations of Theorem 1.2 beyond the matrix case.
The proof of Lemma 1.3, is obtained by retracing the steps in [5]. While the direct repetition of the arguments would not necessarily lead to the exponent $2/5$, one can make certain optimizations along the way, see Subsection 3.2.
Theorem 1.2 is more difficult. While it does not technically involve $C^{*}$-algebras, we believe that this viewpoint best explains our motivation. Note that, if one of the unitaries has a large spectral gap, then the problem becomes “topologically trivial” and, essentially, reduces to the main result of [21], see Proposition 4.1 and Lemma 4.2. Our construction of commuting approximants, which involves several applications of the above “gapped” result, involves a sequence of operator-theoretic procedures (mainly applied to the matrix $u$), and at each step one needs to maintain control over the norm of the commutator with $v$. Not every operator-theoretic procedure respects this property: for example, considering spectral projections or discontinuous functional calculus may blow up the norm of the commutator.
Difficulties of similar kind appeared in the proof of the Brown–Pedersen Conjecture in [23] (see also [24]). In this case, the process could also be described as gap opening by a sequence of operator-theoretic procedures, which have to be performed inside of a particular $C^{*}$-algebra. Again, not all operator-theoretic procedures are $C^{*}$-algebraic”, with the most simple example being the consideration of arbitrary spectral projections.
Our main observation, which appeared in [21] in a somewhat different context, is that many $C^{*}$-algebraic” procedures also have “commutator control” properties. For example, smooth functional calculus for normal elements has this property due to theory of operator Lipschitz functions, see (OL4) and (OL6) in the Appendix. The proof in Section 4.2 is the result of such re-engineering of the whole construction in [23]. While things such as smooth calculus can be considered straightforward, we would like to draw the reader’s attention to the following two much less trivial aspects.
  1. (1)
    The main result of [23] states that one can open a gap in a unitary element of a real rank zero $C^{*}$-algebra $\mathcal{A}$, provided that it belongs to the connected component of the identity in the unitary group. The isospectral homotopy lemma 1.3 is a quantitative commutator analogue of this assumption.
  2. (2)
    The real rank zero property itself states that every self-adjoint element of the $C^{*}$-algebra can be approximated by elements with finite spectra. One can restate it as follows: every self-adjoint element $a\in\mathcal{A}$ is close to another self-adjoint element $a^{\prime}$, such that taking spectral projections of $a^{\prime}$ is a valid $C^{*}$-algebraic operation. The commutator analogue of this property is Claim (4) of Proposition 4.1 (which relies the main results of [21]). Indeed, if $a=a^{*}$ and $[a,v]$ is small, one can apply the claim once and produce an exactly commuting pair $a^{\prime},v^{\prime}]=0$. Afterwards, if $f$ is a bounded Borel function, then $f(a^{\prime}),v^{\prime}]=0$. Since $v$ is close to $v^{\prime}$, the commutator $f(a^{\prime}),v]$ will remain small.
We note that each step of this kind leads to some losses in the commutator norm, and different interpretations of abstract $C^{*}$-algebraic constructions can lead to vastly different quantitative results. Controlling these losses was one of the main technical challenges the proof.
1.3. Structure of the paper
In Section 2, we establish some (mostly well-known) preliminaries from spectral theory, involving estimates of norms of spectral projections associated to disjoint intervals in Subsection 2.1, some general properties of projections and unitary operators in Subsection 2.2, and a version of a more specific result on intertwined families of projections, obtained in [5] in its original form, in Subsection 2.3. In Section 3, we introduce the isospectral invariant in Subsection 3.1, prove Lemma 1.3 (the isospectral homotopy lemma) in Subsection 3.2, and use it to establish equivalence between two invariants in Subsection 3.3. In Section 4, which is the most technical in the paper, we prove Theorem 1.4. Subsection 4.1 contains some preparations, including Lemma 4.2 on almost commuting unitaries with a spectral gap. Following the general ideas in [23], we use Proposition 4.3 to construct an amplification of the original almost commuting pair and perturb it in the larger space to create a spectral gap in Subsection 4.2. This allows to find commuting approximants in a larger space in Subsection 4.3. Afterwards, one needs to descend back into the original space, which is done in two steps, in Subsections 4.4 and 4.5.
2. Preliminaries
In this section, we gather several (mostly already known) results regarding spectral projections, approximate spectral projections, and behavior of spectral projections under small perturbations. The proofs are included mostly for the sake of completeness, as well as in order to introduce notation that will be useful in the later sections.
We will actively use the results from the theory of operator Lipschitz functions, often citing properties (OL1) – (OL8) stated in the Appendix. Let $u_{1},u_{2}\in U(n)$ and $g\colon\mathbb{T}=\mathbb{R}/\mathbb{Z}\to\mathbb{R}$ be a smooth function on the circle. Then $\|g(u_{1})-g(u_{2})\|$ is small provided that $\|u_{1}-u_{2}\|$ is small. As a consequence, if $f$ is another bounded function whose support is disjoint from that of $g$, we have $f(u_{1})g(u_{1})=0$ and therefore $\|f(u_{1})g(u_{2})\|$ is small.
If $g$ is, say, discontinuous, then the first claim from the previous paragraph falls apart. However, some estimates of this kind remain valid if the supports of $f$ and $g$ are sufficiently disjoint. Most of the results below have been used in previous work such as [5, 9], in many cases with better constants. The use of operator Lipschitz functions allows us to provide the proofs in a somewhat unified language. We begin with an elementary fact about bump-type functions.
Proposition 2.1.
Let $\beta>0$ and $I_{1},\ldots,I_{n}\subset\mathbb{T}$ be a system of disjoint intervals on the circle with $\mathrm{dist}(I_{k},I_{\ell})\geq\beta$ for $k\neq\ell$. For $a_{1},\ldots,a_{n}\in[-1,1]$ there exists
$f\in C^{\infty}(\mathbb{T};\mathbb{R}),\quad|f|\leq 1,\quad\left.f\right|_{I_{j}}=a_{j},\quad\|f\|_{\operatorname{OL}(\mathbb{T})}\leq\frac{C}{\beta},$
where $C$ is an absolute constant.
Proof.
From (OL7) in the Appendix, the problem can be restated in the language of $1$-periodic function on $\mathbb{R}$. By rescaling (OL5), at the expense of the factor $1/\beta$ one can consider $\beta=1$. In this case, one can construct $f$ as a linear combination of bump functions with disjoint supports which implies a uniform bound in $C_{b}^{2}(\mathbb{R})$ and allows to complete the proof by applying (OL3). ∎
Remark 2.2.
Note that rescaling must be applied before (OL3), since application to $\|f\|_{C_{b}^{2}(\mathbb{R})}$ directly results in a factor $\beta^{-2}$ instead of $\beta^{-1}$.
2.1. Behavior of spectral projections under small perturbations
The following two lemmas, used in [5, Lemma 2.2 & Lemma 2.9] and other related work, formalize the principle that the orthogonality relations between spectral projections (and, more generally, functions with disjoint supports) are approximately preserved under small perturbations, provided that the boundaries of the intervals under consideration are separated from one another.
Lemma 2.3.
Let $g_{j}\colon\mathbb{T}\to[-1,1]$ be measurable and $u_{j}\in U(n)$, $j=1,2$.
  • $(i)$
    Suppose that $I,J\subset\mathbb{T}$ are disjoint closed intervals and $\mathrm{supp}\,g_{1}\subset I$, $\mathrm{supp}\,g_{2}\subset J$. Then
    $\|g_{1}(u_{1})g_{2}(u_{2})\|\leq\frac{C\|u_{1}-u_{2}\|}{\operatorname{dist}(I,J)}.$
  • $(ii)$
    Suppose that $J\subset I\subset\mathbb{T}$ are intervals and $\mathrm{supp}\,g_{1}\subset J$, $\mathrm{supp}\,(1-g_{2})\subset\mathbb{T}\setminus I$. Then
    $\|g_{2}(u_{2})g_{1}(u_{1})-g_{1}(u_{1})\|\leq\frac{C\|u_{1}-u_{2}\|}{\operatorname{dist}(J,\mathbb{T}\setminus I)}.$
Proof.
In the setting of (i), let $h_{1},h_{2}\colon\mathbb{T}\to[0,1]$ be smooth functions such that
$h_{1}|_{I}\equiv 1,\quad h_{2}|_{J}\equiv 0;\quad h_{1}|_{J}\equiv 0,\quad h_{2}|_{I}\equiv 1,\quad\mathrm{supp}\,h_{1}\cap\mathrm{supp}\,h_{2}=\varnothing.$
From Proposition 2.1, one can choose these functions in a way that $\|h_{j}\|_{\operatorname{OL}(\mathbb{T})}\leq\frac{C}{\operatorname{dist}(I,J)}$, where $C$ is an absolute constant. Then
$\|g_{1}(u_{1})g_{2}(u_{2})\|=\|g_{1}(u_{1})h_{1}(u_{1})h_{2}(u_{2})g_{2}(u_{2})\|\leq\|h_{1}(u_{1})h_{2}(u_{2})\|\\ \leq\|h_{1}(u_{1})h_{2}(u_{1})\|+\|h_{1}(u_{1})\left(h_{2}(u_{2})-h_{2}(u_{1})\right)\|\leq\|h_{2}(u_{2})-h_{2}(u_{1})\|\\ \leq\|h_{2}\|_{\operatorname{OL}(\mathbb{T})}\|u_{1}-u_{2}\|\leq\frac{C\|u_{1}-u_{2}\|}{\operatorname{dist}(I,J)},$
which implies part (i). Here, we used the fact that $h_{1}(u)h_{2}(u)=(h_{1}h_{2})(u)=0$ and the definition of $\|\cdot\|_{\operatorname{OL}}$ (see Appendix). Part (ii) follows from Part (i) applied to the functions $1-g_{2}$ and $g_{1}$. ∎
We denote by $\mathrm{1}_{J}$ the indicator function of a set $J\subset\mathbb{T}$, and by $\mathbf{1}$ the identity operator.
Lemma 2.4.
Let $I,J\subset\mathbb{T}$ be two intervals with disjoint boundaries, and $u_{j}\in U(n)$, $j=1,2$. Then
$\|[\mathrm{1}_{I}(u_{1}),\mathrm{1}_{J}(u_{2})]\|\leq\frac{C\|u_{1}-u_{2}\|}{\operatorname{dist}(\partial I,\partial J)},$
where $C>0$ is an absolute constant.
Proof.
Since $\mathrm{1}_{I}(u_{j})=\mathbf{1}-\mathrm{1}_{\mathbb{T}\setminus I}(u_{j})$, one can replace $I$ by $\mathbb{T}\setminus I$ and/or $J$ by $\mathbb{T}\setminus J$ without loss of generality. As a consequence, the cases where $I$ and $J$ are disjoint, when $I$ and $J$ have disjoint complements, or when one interval is contained in the other follow from Lemma 2.3 (i) (both terms in the commutator can be estimated separately). It remains to consider the case where exactly one endpoint $x$ of $J$ is in the interior of $I$. Let $\beta:=\operatorname{dist}(\partial I,\partial J)$, and consider a smooth bump function $h\colon\mathbb{T}\to[0,1]$ such that $h=1$ in the $\beta/10$-neighborhood of $x$ and $h=0$ outside of the $\beta/5$-neighborhood of $x$. From Proposition 2.1, one can choose $h$ such that $\|h\|_{\operatorname{OL}(\mathbb{T})}\leq C\beta^{-1}$.
By construction, we have $h\leq\mathrm{1}_{I}$, while the function $\mathrm{1}_{I}-h$ vanishes in a neighborhood of $x$. The set $\mathrm{supp}\,(\mathrm{1}_{I}-h)$ is a union of two intervals which we can denote by $I_{1},I_{2}$. Setting $h_{k}:=(1_{I}-h)|_{I_{k}}$ for $k=1,2$, we have
$\mathrm{1}_{I}=h_{1}+h+h_{2},\quad\mathrm{supp}\,h_{j}\subset I_{j},\quad h_{j}\colon\mathbb{T}\to[0,1].$
One of the intervals $I_{j}$ (say, $I_{1}$) must be contained in $J$, and the other (say, $I_{2}$) disjoint from $J$. In both cases, we have $\operatorname{dist}(\partial I_{j},\partial J)\geq\beta/5$. As a consequence,
(2.1) $\|[h_{1}(u_{1}),\mathrm{1}_{J}(u_{2})]\|=\|[h_{1}(u_{1}),\mathrm{1}_{\mathbb{T}\setminus J}(u_{2})]\|\leq 2\|h_{1}(u_{1})\mathrm{1}_{\mathbb{T}\setminus J}(u_{2})\|\leq\frac{10C\|u_{1}-u_{2}\|}{\beta}$
by Lemma 2.3 (i), since $\operatorname{dist}(I_{1},\mathbb{T}\setminus J)\geq\beta/5$. As $\operatorname{dist}(I_{2},\mathbb{T}\setminus J)\geq\beta/5$, we have also
(2.2) $\|[h_{2}(u_{1}),\mathrm{1}_{J}(u_{2})]\|\leq 2\|h_{2}(u_{1})\mathrm{1}_{\mathbb{T}\setminus J}(u_{2})\|\leq\frac{10C\|u_{1}-u_{2}\|}{\beta}.$
Finally, using the smoothness of $h$, we have
(2.3) $\|[h(u_{1}),\mathrm{1}_{J}(u_{2})]\|\leq\|[h(u_{1})-h(u_{2}),\mathrm{1}_{J}(u_{2})\|+\|[h(u_{2}),\mathrm{1}_{J}(u_{2})]\|\leq 2\|h\|_{\operatorname{OL}(\mathbb{T})}\|u_{1}-u_{2}\|\leq\frac{2C\|u_{1}-u_{2}\|}{\beta}.$
Combining the equations (2.1) – (2.3) completes the proof. ∎
2.2. Some general results on orthogonal projections and unitary operators
We now collect several (mostly well-known) results about orthogonal projections, operators close to projections, and almost commuting projections. Recall that every bounded invertible operator $a$ on a Hilbert space admits a polar decomposition
$a=u|a|,$
where $u$ is unitary and $|a|=(a^{*}a)^{1/2}$. In this situation, we will use the notation $\operatorname{Arg}a:=u=a|a|^{-1}$. The following fact is elementary and we omit the proof.
Proposition 2.5.
Let $t_{0}$ be a self-adjoint operator on a Hilbert space such that $\|t_{0}^{2}-t_{0}\|<1/10$. Then the operator
$t:=\frac{1}{2}(\mathbf{1}+\operatorname{Arg}(2t_{0}-\mathbf{1}))$
is an orthogonal projection satisfying
$\|t-t_{0}\|\leq 4\|t_{0}^{2}-t_{0}\|,\quad\operatorname{ran}t\subset\operatorname{ran}t_{0}.$
Remark 2.6.
Note that the assumptions on $t_{0}$ imply that $\sigma(t_{0})\subset[-1/5,1/5]\cup[1-1/5,1+1/5]$. As a consequence, one can also define $t=\mathrm{1}_{[1-1/5,1+1/5]}(t_{0})$.
Remark 2.7.
For an approximate projection operator $t_{0}$ satisfying the assumptions of Proposition 2.5, define
$\operatorname{rank}_{+}(t_{0}):=\operatorname{rank}(t)=\dim\operatorname{ran}\mathrm{1}_{[1-1/5,1+1/5]}(t_{0})$
the dimension of the range of the associated exact projection. Note that $\operatorname{rank}_{+}$ is stable under small perturbations that preserve spectral gap of $t_{0}$ around $1/2$. As a consequence, $\operatorname{rank}_{+}$ is constant on any continuous family of approximate projections satisfying the assumptions of Proposition 2.5.
Suppose also that $\|t_{0}^{2}-t_{0}\|<1/50$, $\|s_{0}^{2}-s_{0}\|<1/50$, $\|t_{0}s_{0}\|<1/50$. Then $t_{0}+s_{0}$ satisfies the assumptions of Proposition 2.5 and
(2.4) $\operatorname{rank}_{+}(t_{0}+s_{0})=\operatorname{rank}_{+}(t_{0})+\operatorname{rank}_{+}(s_{0}).$
We will also need the following version of a well known result about pairs of projections, see [20, Section I.4.6].
Proposition 2.8.
Let $p$ and $q$ be two orthogonal projections on a Hilbert space, with $\|p-q\|<1$. Define
$\sigma:=\left(qp+(\mathbf{1}-q)(\mathbf{1}-p)\right)\left(\mathbf{1}-(p-q)^{2}\right)^{1/2}.$
Then, $\sigma$ is a unitary operator satisfying
$q=\sigma p\sigma^{-1},\quad p=\sigma^{-1}q\sigma.$
Moreover, if $\|p-q\|\leq 1/10$, then
$\|\sigma-\mathbf{1}\|\leq 4\|p-q\|.$
The last estimate can be easily obtained (if needed, with better constants) from the power series expansion of the square root. We will also use the following elementary corollary – again, without optimizing the constants.
Corollary 2.9.
Let $p$ and $q$ be two orthogonal projections on a Hilbert space, with $\|pq\|<1/100$. Then, there exists a unitary operator $\sigma$ such that
$\sigma p\sigma^{-1}\leq\mathbf{1}-q,\quad\|\sigma-\mathbf{1}\|\leq{5}\|pq\|.$
Proof.
Let $t_{0}:=(\mathbf{1}-q)p(\mathbf{1}-q)$. It is easy to see that $t_{0}$ satisfies the assumptions of Proposition 2.5. Denote by $t$ the resulting projection. Note also that the smallness of $\|pq\|$ guarantees that $t_{0}$ (and, therefore, $t$) is close to $p$. Therefore, one can apply Proposition 2.8 to $p$ and $t$. ∎
Remark 2.10.
In Proposition 2.5, Proposition 2.8, and Corollary 2.9, the unitary $\sigma$ is constructed by only using algebraic operations and square roots. As a consequence, both statements hold in any unital $C^{*}$-algebra.
We also formalize the following known claims: an almost unitary operator is close to a unitary operator, compression of a unitary operator by an almost commuting projection is close to a unitary operator, and that large spectral gaps are stable under such compressions. In what follows, $B(z,R)$ denotes an open ball of radius $R$ about $z\in{\mathbb{C}}$. If $q$ is an orthogonal projection and $a$ is an operator, we will use the following shortened notation for the corresponding compression:
$qa|_{q}:=(qaq)|_{\operatorname{ran}q}.$
Proposition 2.11.
  1. $(i)$
    Suppose that $w\in B(H)$ and
    $\|w^{*}w-\mathbf{1}\|\leq\rho<1/5.$
    Then $u=w|w^{*}w|^{-1/2}$ satisfies $\|w-u\|\leq 5\rho$ and $u^{*}u=\mathbf{1}$. In particular, if $\dim H<+\infty$, then $w$ is close to a unitary operator.
  2. $(ii)$
    Suppose that $u\in U(H)$ and $p$ is an orthogonal projection with $\|[p,u]\|\leq\varkappa$. Then there exists a unitary operator $w$ acting on $\operatorname{ran}p$ such that $\|pu|_{p}-w\|\leq 10\varkappa$. Moreover, if $B(z,R)\cap\sigma(u)=\varnothing$, then $B(z,R-10\varkappa)\cap\sigma(w)=\varnothing$.
Proof.
The first claim follows from the spectral mapping theorem applied to $w^{*}w$. In the second claim, existence follows from applying the first claim to $pu|_{p}$. In order to establish the spectral phenomena, note
$\|(w-z)x\|\geq\|(pup-z)x\|-10\varkappa\|x\|=\|(u-z)x\|-10\varkappa\|x\|\geq(R-10\varkappa)\|x\|,\quad\forall x\in\operatorname{ran}p.$
2.3. Unitary equivalence between intertwined families of projections
The following somewhat more specific proposition appears in [5, Lemma 2.8]. We present the statement here with improved estimates (that is, $\sqrt{N}$ instead of $N$ in equation (2.7)) and continued neglect of constant optimization. Since we will need this more explicit formulation, we include a brief argument for the convenience of the reader.
Proposition 2.12.
Let $2\leq N\in\mathbb{N}$, $0\leq{\varepsilon}<1/200$, and $P_{k},Q_{k}$ be a family of orthogonal projections on a Hilbert space satisfying for $j,k\in\mathbb{Z}/N\mathbb{Z}$
$\sum_{k}P_{k}=\sum_{k}Q_{k}=\mathbf{1},$
(2.5) $\|P_{k}Q_{j}-Q_{j}P_{k}\|<{\varepsilon},\quad\|(P_{k}+P_{k+1})Q_{k}-Q_{k}\|<{\varepsilon},\quad\|(Q_{k-1}+Q_{k})P_{k}-P_{k}\|<{\varepsilon},\quad\forall j,k\in\mathbb{Z}/N\mathbb{Z}.$
Then each of the projections $Q_{k}$ and $P_{k}$ admits a decomposition
$P_{k}=p_{2k-1}^{\prime}+p_{2k}^{\prime},\quad Q_{k}=q_{2k}^{\prime}+q_{2k+1}^{\prime}.$
where $p_{j}^{\prime}$ and $q^{\prime}_{j}$ are orthogonal projections for $j\in\mathbb{Z}/2N\mathbb{Z}$ with norm estimates
(2.6) $\|p_{j}^{\prime}-q_{j}^{\prime}\|\leq 200{\varepsilon}.$
Moreover, there is a unitary operator $W$ which satisfies
(2.7) $p^{\prime}_{j}=Wq_{j}^{\prime}W^{-1},\quad j\in\mathbb{Z}/2N\mathbb{Z},\quad\|W-\mathbf{1}\|\leq 100{\varepsilon}\sqrt{N}.$
Proof.
Let
$r_{2k-1}:=P_{k}Q_{k-1}P_{k},\quad r_{2k}:=P_{k}Q_{k}P_{k},\quad s_{2k}:=Q_{k}P_{k}Q_{k},\quad s_{2k+1}:=Q_{k}P_{k+1}Q_{k}.$
$P_{1}$ $Q_{1}$ $P_{1}$ $P_{1}$ $Q_{1}$ $r_{2}=P_{1}Q_{1}P_{1}$
Figure 1.
We visualize these operators according to Figure 1. The estimates of equation (2.5) then imply that each of the quantities
(2.8) $\|r_{j}^{2}-r_{j}\|,\quad\|s_{j}^{2}-s_{j}\|,\quad\|r_{2k-1}r_{2k}\|,\quad\|s_{2k}s_{2k+1}\|,\quad\|r_{j}-s_{j}\|,\quad\|r_{2k-1}+r_{2k}-P_{k}\|,\quad\|s_{2k}+s_{2k+1}-Q_{k}\|$
does not exceed $4{\varepsilon}$ for $k\in\mathbb{Z}/N\mathbb{Z}$, $j\in\mathbb{Z}/2N\mathbb{Z}$. In other words, the families $r_{j}$ and $s_{j}$ can be considered as approximate analogues of the desired $p^{\prime},q^{\prime}$. We will construct the exact versions of $p^{\prime}_{j}$, $q^{\prime}_{j}$ by a series of perturbations of $r_{j}$, $s_{j}$, each of which will be of size $O({\varepsilon})$. We will keep track of explicit constants but will not optimize them.
Let us first apply Proposition 2.5 to each approximate projection $r_{2k-1}$, $r_{2k}$, $s_{2k}$, $s_{2k+1}$, with the Hilbert space being $\operatorname{ran}P_{k}$ or $\operatorname{ran}Q_{k}$, respectively. This yields a family of projections $r_{j}^{\prime},s_{j}^{\prime}$, with the first two quantities in equation (2.8) exactly zero, and the last five quantities not exceeding, say, $16{\varepsilon}$. We also have
$\|r_{j}-r_{j}^{\prime}\|\leq 16{\varepsilon},\quad\|s_{j}-s_{j}^{\prime}\|\leq 16{\varepsilon},\quad\forall j\in\mathbb{Z}/2N\mathbb{Z}.$
We now construct the projections $p_{j}^{\prime},q_{j}^{\prime}$ by applying Corollary 2.9 to each pair $r^{\prime}_{2k-1},r^{\prime}_{2k}$ and $s^{\prime}_{2k},s^{\prime}_{2k+1}$ to conjugate the even indexed projections into the complement of their odd companions inside the range of $P_{k}$ and $Q_{k}$, respectively. This produces unitary operators $U_{k}$, $V_{k}$ acting inside of $\operatorname{ran}P_{k}$, $\operatorname{ran}Q_{k}$ and satisfying
$U_{k}r^{\prime}_{2k}U_{k}^{-1}\leq\mathbf{1}_{P_{k}}-r_{2k-1}^{\prime},\quad V_{k}s^{\prime}_{2k}V_{k}^{-1}\leq\mathbf{1}_{Q_{k}}-s_{2k+1}^{\prime},\quad\|U_{k}-\mathbf{1}_{P_{k}}\|\leq 80{\varepsilon},\quad\|V_{k}-\mathbf{1}_{Q_{k}}\|\leq 80{\varepsilon}.$
Define
$p_{2k-1}^{\prime}:=U_{k}r^{\prime}_{2k-1}U_{k}^{-1},\quad p_{2k}^{\prime}:=P_{k}-p_{2k-1}^{\prime};\quad q_{2k}^{\prime}:=V_{k}s^{\prime}_{2k}V_{k}^{-1},\quad q_{2k+1}^{\prime}:=Q_{k}-q_{2k}^{\prime}.$
We obtain equation (2.6) by combining the above estimates.
It remains to construct the unitary operator $W$. Using equation (2.6), apply Proposition 2.8 to the pair $p^{\prime}_{j},q^{\prime}_{j}$, which will produce a unitary operator $W_{j}$ with
$\|W_{j}-\mathbf{1}\|\leq 400{\varepsilon},\quad W_{j}p_{j}^{\prime}W_{j}^{-1}=q_{j}^{\prime}.$
The operator $W_{j}p_{j}^{\prime}=q_{j}^{\prime}W_{j}$, restricted to the range of $p_{j}^{\prime}$, is a partial isometry between $\operatorname{ran}p_{j}^{\prime}$ and $\operatorname{ran}q_{j}^{\prime}$. Let
$W:=\sum_{j\in\mathbb{Z}/2N\mathbb{Z}}W_{j}p_{j}^{\prime}=\sum_{j\in\mathbb{Z}/2N\mathbb{Z}}q_{j}^{\prime}W_{j}.$
Clearly, $W$ is a unitary operator satisfying $Wp_{j}^{\prime}=q_{j}^{\prime}W$. In order to estimate $\|W-\mathbf{1}\|$, note that for any vector $x$, we have
$\|Wx-x\|^{2}=\sum_{j}\|q_{j}^{\prime}(Wx-x)\|^{2}=\sum_{j}\|W_{j}p^{\prime}_{j}x-q_{j}^{\prime}x\|^{2}\leq\sum_{j}\left(\|(p_{j}^{\prime}-q_{j}^{\prime})x\|^{2}+\|(W_{j}-\mathbf{1})p_{j}^{\prime}x\|\right)^{2}\\ \leq 2\sum_{j}\|(p_{j}^{\prime}-q_{j}^{\prime})x\|^{2}+2\cdot 400^{2}{\varepsilon}^{2}\sum_{j}\|p_{j}^{\prime}x\|^{2}\leq\left(2\cdot 200^{2}N{\varepsilon}^{2}+2\cdot 400^{2}{\varepsilon}^{2}\right)\|x\|^{2}.\,\,\qed$
3. Isospectral invariant, homotopy, and the winding number
In this section we introduce the isospectral invariant $\operatorname{isospec}(u,v)$, originally defined in [5], state our first main result (the quantitative isospectral homotopy lemma), and use it to show that the isospectral invariant coincides with the winding number invariant.
3.1. Definition of the isospectral invariant
Let $I,J\subset\mathbb{T}$ be two intervals whose intersection comprises a single interval. We say that $I$ and $J$ are oriented counterclockwise provided $I\cap J$ prolongs counterclockwise from the (only) point of $\partial J$ which is contained in $I$. For example, if $\omega=e^{2\pi i/8},$ then $\omega,\omega^{3}]$ and $\omega^{2},\omega^{4}]$ are oriented counterclockwise. A version of the following lemma, which serves here as the definition of the isospectral invariant, is established in [5, Theorem 4.1]. However, we need the version with more explicit assumption in equation (3.1) and therefore include the proof.
Proposition 3.1.
There exist an absolute constant $C_{1}>0$ such that for any pair of unitary matrices $u,v\in U(n)$ any two intervals $I,J\subset\mathbb{T}$ oriented counterclockwise such that
(3.1) $1/10\geq\operatorname{dist}(\partial I,\partial J)\geq C_{1}\|[u,v]\|,$
the operator $v\mathrm{1}_{I}(u)v^{*}\mathrm{1}_{J}(u)=\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u)$ is an approximate projection in the sense of Propostion 2.5. The integer number
(3.2) $\operatorname{isospec}(u,v):=\operatorname{rank}(\mathrm{1}_{I}(u)\mathrm{1}_{J}(u))-\operatorname{rank}_{+}(\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u))$
is independent of the choice of $I,J$ within the assumptions of equation (3.1), and remains constant if one continuously varies $u,v$ within the constraints on $\|[u,v]\|$ determined by equation (3.1).
Proof.
Let $C_{1}=100C,$ where $C$ is the universal constant appearing in Lemma 2.4. Equation 3.1 then implies the estimate
$\frac{\|[u,v]\|}{\operatorname{dist}(\partial I,\partial J)}\leq\frac{1}{C_{1}}$
Since $\|u-vuv^{*}\|=\|uv-vu\|$, Lemma 2.4 implies
(3.3) $\|[\mathrm{1}_{I}(vuv^{*}),\mathrm{1}_{J}(u)]\|\leq\frac{C\|uv-vu\|}{\operatorname{dist}(\partial I,\partial J)}\leq\frac{C}{C_{1}}.$
As a consequence, the real part $\Re\left(\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u)\right)$ is an approximate projection in the sense of Proposition 2.5 which is, say, $\frac{1}{100}$-close to a projection for $C_{1}$ large enough. On the other hand, equation (3.3) also implies
$\|\Re\left(\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u)\right)-\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u)\|\leq\frac{1}{100}.$
As a consequence, $\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u)$ is within of $1/50$ norm distance from some orthogonal projection.
In order to establish independence from the choice of intervals, assume that $I=I_{1}\cup I_{2}$ is a union of two non-intersecting intervals such that the pair $I_{1},J$ satisfies equation (3.1) and $I_{2}$ is completely contained inside of $J$ with the same boundary conditions. Then, one has
$\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u)=\mathrm{1}_{I_{1}}(vuv^{*})\mathrm{1}_{J}(u)+\mathrm{1}_{I_{2}}(vuv^{*})\mathrm{1}_{J}(u),$
where each term is $\frac{1}{50}$-close to a projection. As a consequence, $\operatorname{rank}_{+}$ is well-defined and additive, so that
$\operatorname{rank}_{+}(\mathrm{1}_{I}(vuv^{*})\mathrm{1}_{J}(u))=\operatorname{rank}_{+}(\mathrm{1}_{I_{1}}(vuv^{*})\mathrm{1}_{J}(u))+\operatorname{rank}_{+}(\mathrm{1}_{I_{2}}(vuv^{*})\mathrm{1}_{J}(u)).$
On the other hand, Lemma 2.3 (ii) implies that
$\|\mathrm{1}_{I_{2}}(vuv^{*})\mathrm{1}_{J}(u)-\mathrm{1}_{I_{2}}(vuv^{*})\|\leq\frac{C\|u-vuv^{*}\|}{C_{2}\|uv-vu\|}=\frac{C}{C_{1}}<1.$
As a consequence of Proposition 2.8,
$\operatorname{rank}_{+}(\mathrm{1}_{I_{2}}(vuv^{*})\mathrm{1}_{J}(u))=\operatorname{rank}(\mathrm{1}_{I_{2}}(vuv^{*}))=\operatorname{rank}(\mathrm{1}_{I_{2}}(u))=\operatorname{rank}(\mathrm{1}_{I_{2}}(u)\mathrm{1}_{J}(u)),$
which implies that the additional terms in equation (3.2) produced by $I_{2}$ cancel one another. The same arguments apply if $I_{2}$ is completely contained in $\mathbb{T}\setminus J$, and for modifications of the interval $J$ of similar kind. It is easy to see (using the fact that $\operatorname{dist}(\partial I,\partial J)\leq 1/10$) that any pair of counterclockwise oriented intervals satisfying equation (3.1) can be transformed into any other pair by a series of, say, at most 10 such operators (see below). As a consequence, $\operatorname{isospec}(u,v)$ does not depend on the choice of intervals for this choice of $C_{1}$.
For the convenience of the reader, we provide more details about the above transformation. Within three moves one pair $I_{1},J_{1}$ can be arranged so that $I_{1}\cup J_{1}$ is as short as possible; this choice can be made while fixing one endpoint of $I_{1}\cap J_{1}.$ Following the same principle, choose a preferred short arrangement for the target positively oriented pair $I_{2},J_{2}$. Within four moves, the short arrangement we have can be transformed to the preferred short arrangement of the target pair. By reversing the moves which brings the preferred short arrangement to the target pair, we move the given pair to the target pair in ten moves. A more careful inspection of the argument may reduce the number of transformations, which will result in a small improvement in some constants.
It remains to establish homotopy invariance of the left hand side of equation (3.2). Suppose that $u(t),v(t)$ depend continuously on the parameter $t$. Since the spectrum of $u$ consists of finitely many eigenvalues that vary continuously with $t$, for every specific value of $t$ one can modify $I$, $J$ in a way that will not change the right hand side of the equation (3.2), but $\partial I\cup\partial J$ will contain no eigenvalues of $u$ (this modification depends on $t$, not necessarily in a continuous way). Therefore, all four projections involved in the definition of equation (3.2) will be continuous at that value of $t$, which will imply continuity of the associated approximate projections and therefore their ranks. As a consequence, the left hand side of (3.2) is continuous and therefore constant in $t$. ∎
As an example, we calculate the isospectral invariant of the Voiculescu’s unitaries [25]. We note
Proposition 3.2.
For $m=2,3,\ldots$, let $\omega:=e^{2\pi i/m}$ and
$\Omega_{m}:=\begin{bmatrix}\omega&&&\\ &\omega^{2}&&\\ &&\ddots&&\\ &&&\omega^{m}\end{bmatrix}\qquad S_{k}:=\begin{bmatrix}0&1&&\\ &0&\ddots&\\ &&\ddots&1\\ 1&&&0\end{bmatrix}$
Then, the right hand side of (3.2) is equal to 1-1. As a consequence, $\operatorname{isospec}(\Omega_{m},S_{m})=-1$ whenever it is well defined (say, for $m\geq 7$).
Proof.
A direct calculation shows (see, for example, [11]) that for $m\geq 7$
$\|[S_{m},\Omega_{m}]\|\leq|1-e^{2\pi i/m}|\leq\frac{10}{m}.$
As a consequence, the condition in equation (3.1) is satisfied for $m$ large enough. Note, however, that all four indicator functions, involved in equation (3.2), commute with each other, and therefore equation (3.2) is well defined even without assuming equation (3.1). Let $e_{1},\ldots,e_{m}$ be the standard basis in ${\mathbb{C}}^{m}$, and let $p_{j}:=\langle e_{j},\cdot\rangle e_{j}$ be the associated projections. If
(3.4) $I=[\omega^{s_{1}},\omega^{t_{1}}],\quad J=[\omega^{s_{2}},\omega^{t_{2}}],\quad 1\leq s_{1}\leq s_{2}\leq t_{1}\leq t_{2}\leq m,$
then one can easily check
$\mathrm{1}_{I}(\Omega_{m})=p_{s_{1}}+\ldots+p_{t_{1}},\quad\mathrm{1}_{J}(\Omega_{m})=p_{s_{2}}+\ldots+p_{t_{2}},\quad\mathrm{1}_{I}(S_{m}\Omega_{k}S_{m}^{*})=p_{s_{1}-1}+\ldots+p_{t_{1}-1},$
where $p_{0}:=p_{m}$, and therefore the right hand side of equation (3.2) is always equal to 1-1. We note that, assuming that $m$ is large enough, one can always find a pair of intervals among those in equation (3.4) satisfying equation (3.1). However, the reader can also check that the above calculation can be extended for any pair of intervals oriented counterclockwise, regardless of the norm of the commutator or distance estimates, assuming that the union of these intervals does not contain at least one diagonal entry of $\Omega_{m}$. ∎
Remark 3.3.
Proposition 3.1 serves as the definition of $\operatorname{isospec}(u,v)$. Since it relies on existence of at least one choice of intervals $I,J$ satisfying equation (3.1), we have that $\operatorname{isospec}(u,v)$ is only defined under the assumption $\|[u,v]\|<\frac{1}{10C_{1}}$. Rather than imposing this assumption each time $\operatorname{isospec}(u,v)$ is used, we will state the results that follow in a way that makes them trivial if $\|[u,v]\|$ is too large for $\operatorname{isospec}(u,v)$ to be defined.
3.2. Proof of Lemma 1.3
For two unitaries $u$ and $v$ with $\operatorname{isospec}(u,v)=0$, the aim of Lemma 1.3 is to construct a piecewise smooth path $v_{t}\colon[0,1]\to U(n)$ satisfying
$v_{0}=v,\quad v_{1}=\mathbf{1},\quad\|[v(t),u]\|\leq C\|[u,v]\|^{2/5}.$
Let $u,v\in U(n)$ and $\delta:=\|[u,v]\|$. Fix $2\leq N\in\mathbb{N}$, and let
$I_{j}:=\exp\left(\frac{i}{2N}[j,j+1)\right);\quad\lambda_{j}:=\exp\left(\frac{ij}{2N}\right),\quad j\in\mathbb{Z}$
be a system of intervals on the circle, with $\lambda_{j}$ being the endpoint of the corresponding interval. Normally, one would only consider the range $j=0,\ldots,2N-1$, but we include the possibility of arbitrary integer $j$ to take advantage of the periodicity of the exponential function. Let also
(3.5) $q_{j}:=\mathrm{1}_{I_{j}}(u),\quad Q_{k}:=q_{2k}+q_{2k+1},\quad\tilde{u}:=\sum_{j=0}^{2N-1}\lambda_{j}q_{j}.$
Clearly, we have
$\|\tilde{u}-u\|\leq\frac{2\pi}{N}.$
Moreover, supposing that $z$ is a unitary operator commuting with all $Q_{k}$, $k=0,\ldots,N$, it follows that
(3.6) $\|z\tilde{u}z^{*}-\tilde{u}\|\leq\frac{8\pi}{N}.$
Indeed, since both $\tilde{u}$ and $z$ commute with each $Q_{k}$, it is sufficient to consider restrictions of both operators into the range of $Q_{k}$. However, such restriction of $\tilde{u}$ is a perturbation of a multiple of the identity of size at most $\frac{4\pi}{N}$. Any unitary conjugation by $z$ will be an operator of the same kind. Similarly to equation (3.5), define
$p_{j}:=\mathrm{1}_{I_{j}}(vuv^{*})=vq_{j}v^{*},\quad P_{k}:=p_{2k-1}+p_{2k}.$
The estimates in equation (3.6) also hold for a unitary defined from linear combinations of $p_{k}.$
Assume that $C\delta N<1/200$, where $C$ is the constant from Lemma 2.4 (which exceeds that from Lemma 2.3). Then, one can easily check using Lemmas 2.3 and 2.4 that the families $P_{k}$, $Q_{k}$ satisfy the assumptions of Proposition 2.12,
$\|[P_{j},Q_{k}]\|\leq C\delta N\quad\|(P_{j}+P_{j+1})Q_{j}\|\leq C\delta N\quad\|(Q_{k-1}+Q_{k})P_{k}\|\leq C\delta N.$
This produces another partition
$P_{k}=p_{2k-1}^{\prime}+p_{2k}^{\prime},\quad Q_{k}=q_{2k}^{\prime}+q_{2k+1}^{\prime},\quad k=0,\ldots,2N-1.$
Following the periodicity conventions, we have $p_{-1}=p_{2N-1}$. Moreover, each pair of intervals $I=I_{2k-1}\cup I_{2k}$, $J=I_{2k}\cup I_{2k+1}$ satisfies the assumptions of Proposition 3.1, and, appealing to (2.7) in Proposition 2.12, it follows that
(3.7) $0=\operatorname{isospec}(u,v)=\operatorname{rank}p_{j}^{\prime}-\operatorname{rank}p_{j}=\operatorname{rank}q_{j}^{\prime}-\operatorname{rank}q_{j},\quad\forall j=0,\ldots,2N-1,$
assuming, say, $C_{1}\delta N<1/100$, where $C_{1}$ is the constant from Proposition 3.1. Since $p_{j}$ and $q_{j}$ are also conjugated by $v$, we have that all four projections $p_{j},q_{j},p_{j}^{\prime},q_{j}^{\prime}$ are of equal rank. Our next step shows that in fact there is also a proximity relation between $p_{j}$ and $p_{j}^{\prime}$, and between $q_{j}$ and $q_{j}^{\prime}$.
We will now introduce several unitary operators involved in the construction of the path.
  1. (1)
    For $k=0,\ldots,N-1$, denote by $z_{k}$ any unitary operator acting inside of $\operatorname{ran}Q_{k}$ satisfying
    $q_{2k}^{\prime}=z_{k}q_{2k}z_{k},\quad z_{k}q_{2k+1}z_{k}^{*}=q_{2k+1}^{\prime}.$
    Note that the second property follows from the first one since $Q_{k}=q_{2k}+q_{2k+1}=q_{2k}^{\prime}+q_{2k+1}^{\prime}$. The existence of such $z_{k}$ follows from equation (3.7).
  2. (2)
    Let $z:=\oplus_{k}z_{k}$ be the combined unitary operator. There exists a self-adjoint operator $h$ with $\|h\|\leq 2\pi$ such that $z=e^{ih}$. For $t\in[0,1]$, let $z_{t}:=e^{ith}$.
  3. (3)
    Let $w$ be the operator constructed during the earlier application of Proposition 2.12, satisfying $\|w-\mathbf{1}\|<100\varepsilon\sqrt{N}$, where ${\varepsilon}=CN\delta$. Denote by $w_{t}$ the shortest unitary path such that $w_{0}=\mathbf{1}$, $w_{1}=w$.
  4. (4)
    Similarly to the first part, for $k=0,\ldots,N-1$, denote by $y_{k}$ any unitary operator acting inside of $\operatorname{ran}P_{k}$ satisfying
    $p_{2k}^{\prime}=y_{k}p_{2k}y_{k}^{*},\quad y_{k}p_{2k+1}y_{k}^{*}=p_{2k+1}^{\prime}.$
  5. (5)
    Finally, let $y:=\oplus_{k}y_{k}=e^{ig}$, $g=g^{*}$ $\|g\|\leq 2\pi$, and $y_{t}:=e^{itg}$.
With the above preparations, consider the following path $\{\Gamma_{t}\colon t\in[0,3]\}$ of unitary operators,
(3.8) $\Gamma_{t}:=\begin{cases}z_{t},&t\in[0,1)\\ w_{t-1}z&t\in[1,2)\\ y_{t-2}wz&t\in[2,3].\end{cases}$
Equation (3.6) and the estimate $\|w-\mathbf{1}\|<100\varepsilon\sqrt{N}=100CN^{3/2}\delta$ imply
(3.9) $\Gamma_{0}\tilde{u}\Gamma_{0}^{*}=\tilde{u},\quad\Gamma_{3}\tilde{u}\Gamma_{3}^{*}=\sum_{j=0}^{2N-1}\lambda_{j}p_{j}=v\tilde{u}v^{*},\quad\|\Gamma_{t}\tilde{u}\Gamma_{t}^{*}-\tilde{u}\|\leq C\left(N^{3/2}\delta+\frac{1}{N}\right),\quad t\in[0,3],$
where $C$ is an absolute constant.
Let
$v_{t}:=\Gamma_{t}^{*}v,\quad t\in[0,3].$
Then
$v_{3},\tilde{u}]=0,\quad\|v_{t}\tilde{u}v_{t}^{*}-\tilde{u}\|=\|v\tilde{u}v^{*}-\Gamma_{t}\tilde{u}\Gamma_{t}^{*}\|\leq\|\Gamma_{t}\tilde{u}\Gamma_{t}^{*}-\tilde{u}\|+\|\tilde{u}-v\tilde{u}v^{*}\|\leq C\left(N^{3/2}\delta+\frac{1}{N}\right).$
We have connected $v_{0}=v$ with $v_{3}$ which commutes with $\tilde{u}$. It remains to connect $v_{3}$ to $\mathbf{1}$ in a way that continues to commute with $\tilde{u}$. To do so, simply consider the restriction of $v_{3}$ into an eigenspace of $\tilde{u}$ and consider any unitary path inside that eigenspace that connects this restriction to the identity. For $t\in[3,4]$, denote by $v_{t}$ the resulting family of unitary operators, produced by performing that step in each eigenspace. The choice $N^{-1}=N^{3/2}\delta$, in order to balance the bounds, ultimately leads to
$\|[v_{t},\tilde{u}]\|\leq C\delta^{2/5},\quad\|[v_{t},u]\|\leq\|[v_{t},\tilde{u}]\|+\|\tilde{u}-u\|\leq(C+4\pi)\delta^{2/5},\quad t\in[0,4].\,\,\qed$
3.3. Equivalence between the invariants
We now establish the equivalence of the isospectral invariant with the winding number invariant. This equivalence appears in the literature [11, 5] by identifying each obstruction with the Bott element. However, a more direct proof is possible following the same steps as [5, Section 9], without involving the Bott invariant, as the following quantitative corollary of Lemma 1.3.
Corollary 3.4.
Let $u,v\in U(n)$ with $C\|[u,v]\|^{2/5}<1$, where $C$ is the constant from Lemma 1.3. Define $w(u,v)$ to be the winding number of the curve
$\omega\colon[0,1]\to\mathbb{C}\setminus\{0\},\quad\omega(t):=\det(t\cdot uv+(1-t)\cdot vu).$
Then $w(u,v)=\operatorname{isospec}(u,v)$.
Proof.
We note that that $w(\cdot,\cdot)$ is a homotopy invariant within the constraint $\|[u,v]\|<1$. As a consequence, under the assumptions of Lemma 1.3 with $C\delta^{2/5}<1$, we have $w(u_{t},v)=\mathrm{const}$ along the path provided by that lemma. As a consequence,
$w(u,v)=w(u_{0},v)=w(u_{1},v)=w(\mathbf{1},v)=0,\quad\text{whenever}\quad\operatorname{isospec}(u,v)=0.$
In order to treat the remaining cases, we will make use of Voiculescu’s unitaries $S_{m},\Omega_{m}\in U(m)$, introduced in Proposition 3.2; the latter, combined with [11], implies
$\|[S_{m},\Omega_{m}]\|\leq|1-e^{2\pi i/m}|\leq\frac{10}{m};\quad w(S_{m},\Omega_{m})=\operatorname{isospec}(S_{m},\Omega_{m})=-1,\quad m\geq 7.$
Both $\operatorname{isospec}(\cdot,\cdot)$ and $w(\cdot,\cdot)$ are additive with respect to direct sums:
$\operatorname{isospec}(u_{1}\oplus u_{2},v_{1}\oplus v_{2})=\operatorname{isospec}(u_{1},v_{1})+\operatorname{isospec}(u_{2},v_{2});\quad w(u_{1}\oplus u_{2},v_{1}\oplus v_{2})=w(u_{1},v_{1})+w(u_{2},v_{2}).$
Suppose now that $u,v$ satisfy the assumptions of the corollary and $\operatorname{isospec}(u,v)>0$. Let
$U:=u\oplus u^{\prime},\quad V=v\oplus v^{\prime},$
where $u^{\prime},v^{\prime}$ are made of $\operatorname{isospec}(u,v)$ copies of $\Omega_{m},S_{m}$ with $m$ large enough so that one still has $C\|[U,V]\|^{2/5}<1$. Due to additivity, one has $\operatorname{isospec}(U,V)=0$, which implies, as in the beginning of the proof, $w(U,V)=0$, and therefore
$w(u,v)=-w(u^{\prime},v^{\prime})=-\operatorname{isospec}(u^{\prime},v^{\prime})=\operatorname{isospec}(u,v).$
In order to deal with $\operatorname{isospec}(u,v)<0$, one can replace $S_{m},V_{m}$ by $V_{m}^{-1}$, $S_{m}^{-1}$, respectively, which will reverse the signs of both invariants, and apply the same arguments. ∎
Remark 3.5.
If $u^{\prime},v^{\prime}]=0$, then, clearly $w(u^{\prime},v^{\prime})=\operatorname{isospec}(u^{\prime},v^{\prime})=0$. As a consequence,
$w(u,v)=\operatorname{isospec}(u,v)=0\quad\text{whenever}\quad\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq\frac{1}{10}.$
We arrive to the well-known fact that the assumption $w(u,v)=0$ is indeed necessary in Theorem 1.2.
4. Commuting approximants and topological triviality
In this section we prove Theorem 1.4. We will provide a mechanism of how the outcome of Lemma 1.3 allows to reduce the version of Question 1.1 in Theorem 1.2 to another question, which is “topologically equivalent” to Lin’s theorem, where in each step we keep track of quantitative estimates.
4.1. Some preparations
The following proposition combines improved versions of several well-known results, see, for example, see, for example [18, Section 2]. The original proof in [18] relied on the outcome of [17] as a “black box”. In the following proof, which is included mainly for the convenience of the reader, we replace the use of [17] by the stronger result of [21].
Proposition 4.1.
Suppose that $t,s\in M_{n}(\mathbb{C})$ satisfy one of the following properties:
  • $t=t^{*}$ and $s=s^{*}$;
  • $s=t^{*}$;
  • $s=t^{*}$ and $\mathbf{1}\leq|t|\leq 3\cdot\mathbf{1}$;
  • $t=t^{*}$, $\|t\|\leq 1$, and $s\in U(n)$.
Then there exist matrices $s^{\prime},t^{\prime}$ satisfying the same respective conditions such that
$t^{\prime},s^{\prime}]=0\quad\|t-t^{\prime}\|+\|s-s^{\prime}\|\leq C\big\|[t,s]\big\|^{1/2},\quad\|t^{\prime}\|\leq\|t\|,\quad\|s^{\prime}\|\leq\|s\|.$
Proof.
The first claim is the original form of Lin’s theorem, in the quantitative version from [21]. As explained earlier in the Introduction, the second claim follows from the first one applied to the real and imaginary parts of $s$. Note that in the first two claims we do not assume $s$ and $t$ to be contractions, since the exponent $1/2$ makes the problem scale-invariant. The matrices $s^{\prime}$ and $t^{\prime}$ can be normalized in order to satisfy the norm requirement, which may result in a modification of $C$.
In order to obtain the third claim, start from applying the second claim and assume (without loss of generality, perhaps after a modification of $C$) that $\|t-t^{\prime}\|=\|s-s^{\prime}\|\leq 1/10$. Since the function $z\mapsto|z|$ is operator Lipschitz on any compact subset of $\mathbb{C}\setminus\{0\}$, we have that
$\||s|-|s^{\prime}|\|=\||t|-|t^{\prime}|\|\leq C^{\prime}\|[t,s]\|^{1/2}\leq 1/10,$
where, again, the last inequality can also be assumed without loss of generality by choosing $C$ large enough. As a consequence,
$\left(1-C^{\prime}\|[t,s]\|^{1/2}\right)\mathbf{1}\leq|t^{\prime}|\leq 3\cdot\mathbf{1}.$
In order to satisfy the second part of the third claim, let
$g(z):=\begin{cases}\frac{z}{|z|},&|z|\leq 1,\\ z,&1<|z|<3,\\ 3\frac{z}{|z|}&|z|\geq 3\end{cases}$
defined on $\mathbb{C}\setminus\{0\}$. Since both $s^{\prime}$ and $t^{\prime}$ are normal, the standard functional calculus (spectral mapping theorem) implies $1\cdot\mathbf{1}\leq|g(t^{\prime})|\leq 3\cdot\mathbf{1}$, as well as
$\|s-g(s^{\prime})\|+\|t-g(t^{\prime})\|\leq\|s-s^{\prime}\|+\|s^{\prime}-g(s^{\prime})\|+\|t-t^{\prime}\|+\|t^{\prime}-g(t^{\prime})\|\leq C^{\prime\prime}\|[t,s]\|^{1/2}.$
In order to establish the fourth claim, let $a:=s(t+2\cdot\mathbf{1})$. Since $\mathbf{1}\leq t+2\cdot\mathbf{1}\leq 3\cdot\mathbf{1}$, we have that $a$ and $a^{*}$ satisfy the assumptions of the third claim with, say,
$\|[a,a^{*}]\|\leq 10\|[s,t]\|^{1/2}.$
Let $b$ be a normal operator obtained by applying the third claim, and
$t^{\prime}:=|b|-2=(b^{*}b)^{1/2}-2,\quad s^{\prime}:=b|b|^{-1}=b(b^{*}b)^{-1/2}.$
We have
$t-t^{\prime}=(b^{*}b)^{1/2}-(a^{*}a)^{1/2},\quad s-s^{\prime}=(b-a)(b^{*}b)^{-1/2}+a\left((b^{*}b)^{-1/2}-(a^{*}a)^{-1/2}\right).$
Now, the claim follows from the estimates
$\|(b^{*}b)^{1/2}-(a^{*}a)^{1/2}\|+\|(b^{*}b)^{-1/2}-(a^{*}a)^{-1/2}\|\leq C\|b^{*}b-a^{*}a\|\leq 20C\|b-a\|\leq C^{\prime}\|[s,t]\|,$
in which we used the facts that
$\mathbf{1}\leq b^{*}b\leq 10\cdot\mathbf{1},\quad\mathbf{1}\leq a^{*}a\leq 10\cdot\mathbf{1},$
and that both functions $x\mapsto x^{1/2}$ and $x\mapsto x^{-1/2}$ are operator Lipschitz on $1,10]\subset\mathbb{R}$. ∎
It is not hard to show that if one replaces $t$ in the last part of Proposition 4.1 by a unitary operator with a spectral gap, the approach would still work. However, the size of the gap will appear in the final estimate. The following lemma quantifies this observation. We note that while it is natural to expect the extra factor $\varepsilon^{-1/2}$, where $\varepsilon$ is the size of the gap, it requires a somewhat delicate estimate of the operator Lipschitz norm of a branch of the argument function, which does not offer a natural rescaling property. This estimate is provided by Lemma A.1 in the Appendix. Modulo that estimate, the proof is very short.
Lemma 4.2.
Suppose, $u$ and $v$ are two unitary matrices. Assume, in addition, that $u$ has a spectral gap of size $\rho>0$. Then there exist unitary $u^{\prime},v^{\prime}$ such that
(4.1) $u^{\prime},v^{\prime}]=0,\quad\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq C\rho^{-1/2}\|[u,v]\|^{1/2}.$
Proof.
Without loss of generality, one can assume that the spectral gap of $u$ is around 1-1. Let $x:=\frac{1}{2\pi}\arg u=\frac{1}{2\pi}\arg_{\rho}u$ as considered in the Appendix. From Lemma A.1 and (OL6), we have
$\|[x,v]\|\leq C\rho^{-1}\|[u,v]\|.$
Apply the last claim of Proposition 4.1 with $t=x$ and $s=v$, thus arriving to a commuting pair of $x^{\prime}$ and $v^{\prime}$. Since the function $x\mapsto e^{ix}$ is operator Lipschitz on $\mathbb{R}$, we have that $u^{\prime}=e^{2\pi ix^{\prime}}$ and $v^{\prime}$ satisfy equation (4.1). ∎
Lemma 4.2 provides a strategy for proving Theorem 1.2 if one can answer the following question: given two unitary matrices $u,v$ with small commutator, is it possible to “open a gap” in the spectrum of $u$ by a small perturbation, while preserving smallness of the commutator with $v$ in an appropriate (uniform in the dimension and independent of the perturbation size) sense? In view of Remark 3.5, a positive answer to this question must rely on vanishing of the winding number $w(u,v)$. The following result, established (with a short proof) in [24], shows that such gap opening is possible, regardless of the value of $w(u,v)$, if one considers an amplification by $u^{*}$.
Proposition 4.3.
Let $\mathcal{A}$ be a unital $C^{*}$-algebra, and $u\in U(\mathcal{A})$ be a unitary element. For ${\varepsilon}>0$, let
(4.2) $w({\varepsilon}):=\begin{pmatrix}u&0\\ 0&I\end{pmatrix}\begin{pmatrix}\cos(\pi/2-{\varepsilon})&\sin(\pi/2-{\varepsilon})\\ -\sin(\pi/2-{\varepsilon})&\cos(\pi/2-{\varepsilon})\end{pmatrix}\begin{pmatrix}u^{*}&0\\ 0&I\end{pmatrix}\begin{pmatrix}\cos(\pi/2-{\varepsilon})&-\sin(\pi/2-{\varepsilon})\\ \sin(\pi/2-{\varepsilon})&\cos(\pi/2-{\varepsilon})\end{pmatrix}$.
Then, for $0<{\varepsilon}<1/10$, we have:
$\left\|w({\varepsilon})-\begin{pmatrix}u&0\\ 0&u^{*}\end{pmatrix}\right\|<3{\varepsilon},\quad\|(w(\varepsilon)+\mathbf{1})^{-1}\|\leq{\varepsilon}^{-1}.$
Remark 4.4.
From equation (4.2), one can easily observe
$\|\left[w(\varepsilon),\mathrm{diag}(v,v)\right]\|\leq 2\|[u,v]\|,\quad\forall v\in\mathcal{A}.$
Note that $w(0)=\text{diag}(u,u^{*})$. Therefore, equation (4.2) provides a perturbation of $\text{diag}(u,u^{*})$ of size $O({\varepsilon})$ that opens a gap of size $\varepsilon$} (the extra factors are mostly to absorb the difference between arc length and diameter). As mentioned earlier, there are no additional assumptions on $\mathcal{A}$ and $u$. In principle, any topological obstruction (such as the winding number invariant), preventing the opening of a gap in the spectrum of $u$, exactly cancels with that of $u^{*}$.
4.2. Proof of Theorem 1.4: the amplified almost commuting pair with a spectral gap
As a reminder, we assume that $u,v$ are unitary matrices such that $\|[u,v]\|\leq\delta$ and there exists a continuous path $\{u_{t}\colon t\in[0,1]\}$ satisfying
$u_{0}=u,\quad u_{1}=\mathbf{1},\quad\|[u_{t},v]\|\leq\delta\quad\forall t\in[0,1].$
To prove Theorem 1.4 we must construct unitary matrices $u^{\prime}$, $v^{\prime}$ such that
$u^{\prime},v^{\prime}]=0,\quad\|u-u^{\prime}\|+\|v-v^{\prime}\|\leq C\delta^{1/12}.$
Note that, in view of Lemma 4.2, it would be sufficient to transform $u$ and $v$ into a pair of almost commuting unitary matrices where, in addition, one of them has a spectral gap. Following the strategy of retracing the steps of [23], we first consider amplifications of the original pair $(u,v)$ in some space of larger dimension, in which creating a gap would be easier. Afterwards, we carefully descend back into the original space in order to produce a commuting pair. The descent will be performed in two steps, with each step being referred to as “dimension reduction”.
Throughout, we denote by $A$ the matrix algebra over an arbitrary finite dimensional vector space – the exact dimension will not influence the result. So, for $k\in\mathbb{N}$, denote by $A_{k}=M_{k}(A)=A\otimes M_{k}(\mathbb{C})$ the $k$th matrix amplification of $A.$
Now, fix $\varepsilon>0.$ Under the assumptions of Theorem 1.4, choose a (large) integer $d\in\mathbb{N}$ and subdivide the path $u_{t}$ into $d$ segments of length at most $\varepsilon$}; that is, construct a sequence of matrices $u_{0},\ldots,u_{d}$ such that
$u_{0}=u,\quad u_{d}=\mathbf{1},\quad\|[u_{j},v]\|\leq\delta,\quad\|u_{j+1}-u_{j}\|\leq{\varepsilon}\,\,\,\text{for}\,\,\,j=0,\ldots d-1.$
Along the lines of [23], let
$u_{\mathrm{path}}:=\mathrm{diag}\{u_{1}^{*},u_{1},u_{2}^{*},u_{2},\ldots,u_{d-1}^{*},u_{d-1},\mathbf{1}_{A}\}\in A_{2d-1};$
$u_{\mathrm{amp}}:=\mathrm{diag}\{u,u^{*},u_{1},u_{1}^{*},u_{2},u_{2}^{*},\ldots,u_{d-1},u_{d-1}^{*}\}\in A_{2d}.$
From the construction, it follows
$\|u\oplus u_{\mathrm{path}}-u_{\mathrm{amp}}\|\leq{\varepsilon};\quad\|[u_{\mathrm{path}},v_{\mathrm{path}}]\|\leq\delta,\quad\|[u_{\mathrm{amp}},v_{\mathrm{amp}}]\|\leq\delta,$
where
$v_{\mathrm{path}}:=v\otimes\mathbf{1}_{(2d-1)\times(2d-1)}\in A_{2d-1},\quad v_{\mathrm{amp}}:=v\otimes\mathbf{1}_{2d\times 2d}=v\oplus v_{\mathrm{path}}\in A_{2d}$
are the amplifications of $v$ of the corresponding dimensions.
If one ignores the identity matrix in $u_{\mathrm{path}}$, both $u_{\mathrm{amp}}$ and $u_{\mathrm{path}}$ are made out of blocks of the form that allows the application of Proposition 4.3. For $0<{\varepsilon}<1/10$, denote the results of applying that proposition inside each block by $u_{\mathrm{path}}({\varepsilon})$ and $u_{\mathrm{amp}}({\varepsilon})$, respectively, without altering the last block entry $\mathbf{1}$ in $u_{\mathrm{path}}$. It is easy to see that they satisfy
$\|u_{\mathrm{path}}({\varepsilon})-u_{\mathrm{path}}\|\leq 3{\varepsilon},\quad\|u_{\mathrm{amp}}({\varepsilon})-u_{\mathrm{amp}}\|\leq 3{\varepsilon};\quad\|[u_{\mathrm{path}}({\varepsilon}),v_{\mathrm{path}}]\|\leq 2\delta,\quad\|[u_{\mathrm{amp}}({\varepsilon}),v_{\mathrm{amp}}]\|\leq 2\delta.$
The last two estimates follow from the precise forms of $u_{\mathrm{path}}(\varepsilon)$ and $u_{\mathrm{amp}}(\varepsilon)$ in Proposition 4.3 on each $2\times 2$ block. Moreover, both $u_{\mathrm{path}}({\varepsilon})$ and $u_{\mathrm{amp}}({\varepsilon})$ have spectral gaps of size $\varepsilon$} near 1-1.
It also follows that
$\|u\oplus u_{\mathrm{path}}({\varepsilon})-u_{\mathrm{amp}}({\varepsilon})\|\leq\|u\oplus u_{\mathrm{path}}-u_{\mathrm{amp}}\|+6\varepsilon\leq 7\varepsilon.$
If $p$ is the projection onto the copy of $A_{d}$ associated to the top left corner, we have, in the above notation,
$\|u-pu_{\mathrm{amp}}(\varepsilon)|_{p}\|=\|p(u_{\mathrm{amp}}-u_{\mathrm{amp}}(\varepsilon))p\|\leq 3\varepsilon.$
As a consequence, one can consider $u_{\mathrm{amp}}(\varepsilon)$ as an amplified approximant to $u$, which has a spectral gap of size approximately $\varepsilon$, but also retains approximate commutation relation with $v_{\mathrm{amp}}$ of magnitude approximately $\delta$, where $\varepsilon$ and $\delta$ are not bound by any additional relation.
4.3. Proof of Theorem 1.4: the amplified commuting pair.
With the assurance of a unitary $u_{\mathrm{amp}}(\varepsilon)$ admitting a gap in its spectrum, we apply Lemma 4.2 with $\rho=\varepsilon$ to $u_{\mathrm{amp}}(\varepsilon)$ and $v_{\mathrm{amp}}$ to obtain the commuting pair
$u^{\prime}_{\mathrm{amp}}(\varepsilon),v^{\prime}_{\mathrm{amp}}(\varepsilon)]=0$
subject to
$\|u^{\prime}_{\mathrm{amp}}(\varepsilon)-u_{\mathrm{amp}}(\varepsilon)\|+\|v^{\prime}_{\mathrm{amp}}(\varepsilon)-v_{\mathrm{amp}}\|\leq C\varepsilon^{-1/2}\delta^{1/2}.$
Here, $v_{\mathrm{amp}}^{\prime}(\varepsilon)$ serves as an amplified approximant to $v$, which carries along a commuting unitary $u^{\prime}_{\mathrm{amp}}(\varepsilon).$ We likewise generate the commuting pair $(u^{\prime}_{\mathrm{path}}(\varepsilon),v^{\prime}_{\mathrm{path}}(\varepsilon))$ with identical estimates, from the pair $(u_{\mathrm{path}}(\varepsilon),v_{\mathrm{path}})$. On account of the essential feature of commutation for the unitaries and the estimate
$\|u\oplus u^{\prime}_{\mathrm{path}}(\varepsilon)-u^{\prime}_{\mathrm{amp}}(\varepsilon)\|\leq 7\varepsilon+C\varepsilon^{-1/2}\delta^{1/2},$
we can now abandon our original amplified approximant $u_{\mathrm{amp}}(\varepsilon)$ of $u$ in favor of $u^{\prime}_{\mathrm{amp}}(\varepsilon)$, at the expense of the above estimate.
Since $u^{\prime}_{\mathrm{amp}}(\varepsilon),v^{\prime}_{\mathrm{amp}}(\varepsilon)]=0$, we have $f(u^{\prime}_{\mathrm{amp}}(\varepsilon)),v^{\prime}_{\mathrm{amp}}(\varepsilon)]=0$ for every Borel function $f\colon{\mathbb{C}}\to[0,1]$. As a consequence,
$\left\|[f(u^{\prime}_{\mathrm{amp}}(\varepsilon)),v_{\mathrm{amp}}]\right\|\leq 2\left\|v^{\prime}_{\mathrm{amp}}({\varepsilon})-v_{\mathrm{amp}}\right\|\leq C\varepsilon^{-1/2}\delta^{1/2}.$
In particular, this allows us to create additional gaps in the spectrum of $u^{\prime}_{\mathrm{amp}}(\varepsilon)$ or split into parts using spectral projections, without further losses in the commutator norm. This idea appears several times in the next subsections.
4.4. Proof of Theorem 1.4: the first dimension reduction.
We now briefly summarize our position. Given a pair of unitaries $u,v\in A$, a path $u_{t}$ from $u$ to the identity with $\|[u_{t},v]\|<\delta$, and $\varepsilon<1/10$, we have found integer $d$ and unitaries $u^{\prime}_{\mathrm{amp}}(\varepsilon),v^{\prime}_{\mathrm{amp}}(\varepsilon)\in A_{2d}$ and $u^{\prime}_{\mathrm{path}}(\varepsilon),v^{\prime}_{\mathrm{path}}(\varepsilon)\in A_{2d-1}$ which obey the relations
(4.3) $\|u\oplus u^{\prime}_{\mathrm{path}}({\varepsilon})-u^{\prime}_{\mathrm{amp}}({\varepsilon})\|\leq{\varepsilon}+6{\varepsilon}+C{\varepsilon}^{-1/2}\delta^{1/2},\quad\|[u^{\prime}_{\mathrm{path}}({\varepsilon}),v^{\prime}_{\mathrm{path}}({\varepsilon})]\|=\|[u^{\prime}_{\mathrm{amp}}({\varepsilon}),v^{\prime}_{\mathrm{amp}}({\varepsilon})]\|=0,$
(4.4) $\|v^{\prime}_{\mathrm{path}}({\varepsilon})-v_{\mathrm{path}}\|+\|v^{\prime}_{\mathrm{amp}}({\varepsilon})-v_{\mathrm{amp}}\|\leq C{\varepsilon}^{-1/2}\delta^{1/2}.$
Up to amplification to $A_{2d},$ we have found commuting approximants. The main challenge now is to descend into the original space. Clearly, if we simply apply the compression by the projection associated to the top left corner, the operators will be almost unitary and almost commuting, but we lose control over the spectral gap, in some sense returning to the original problem. A similar difficulty appears in [23]: compression of an element with finite spectrum does not have to be close to an element with finite spectrum. However, compression of a self-adjoint element is still self-adjoint, so it will always be of this kind. The same holds for unitary elements with large gaps. So, one can prepare for the compression by creating such a gap.
We will use spectral projections of $u\oplus u^{\prime}_{\mathrm{path}}(\varepsilon)$ to shed the extra dimensions of the pair $(u^{\prime}_{\mathrm{amp}}(\varepsilon),v^{\prime}_{\mathrm{amp}}(\varepsilon))$ with the aim of drawing the compressions back into commutation through repeated application of Lemma 4.2. To ease the burden of notation, we will proceed denoting the unitaries above by $u^{\prime}_{\mathrm{amp}},v^{\prime}_{\mathrm{amp}},$ and $u^{\prime}_{\mathrm{path}},v^{\prime}_{\mathrm{path}}$, dropping the dependence on $\varepsilon$.
Concerning the operator $u\oplus u^{\prime}_{\mathrm{path}}$, write $p$ for the projection onto the range of $u$ in $A_{2d}$ (that is, the top left corner). Next, we split the spectrum of $u^{\prime}_{\mathrm{path}}$ into two semi-circles: let
$q:={1}_{\{\Re z\leq 0\}}(u^{\prime}_{\mathrm{path}}),\quad r:={1}_{\{\Re z>0\}}(u^{\prime}_{\mathrm{path}});\quad w^{\prime}_{-}:=qu^{\prime}_{\mathrm{path}}|_{q},\quad w^{\prime}_{+}:=ru^{\prime}_{\mathrm{path}}|_{r}$
denote the corresponding spectral projections and components of $u^{\prime}_{\mathrm{path}}$ in the corresponding subspaces, so that
$u\oplus u^{\prime}_{\mathrm{path}}=u\oplus w^{\prime}_{-}\oplus w^{\prime}_{+},\quad\mathbf{1}:=\mathbf{1}_{A_{2d}}=p+q+r.$
Recalling the commentary preceding Lemma 2.3 in Section 2, let $\Omega_{-}=\{\Re{z}\leq-1/2\}\cap\mathbb{T}$ be a closed sub-arc of the closed left semicircle whose boundary is separated from the boundary of $\{\Re z>0\}\cap\mathbb{T}$. Then let $\eta_{-}\colon\mathbb{T}\to[0,1]$ be a smooth function, with $\eta|_{\Omega_{-}}=1$ and vanishing outside a (say) $1/10$-neighborhood of $\Omega_{-}.$
For the corresponding spectral projection $1_{\Omega_{-}}(u^{\prime}_{\mathrm{amp}})$, of $u^{\prime}_{\mathrm{amp}}$, since
$\eta_{-}(u\oplus u^{\prime}_{\mathrm{path}})r=r\eta_{-}(u\oplus u^{\prime}_{\mathrm{path}})=0,$
compute
$\|1_{\Omega_{-}}(u^{\prime}_{\mathrm{amp}})r\|\leq\|\eta_{-}(u^{\prime}_{\mathrm{amp}})r\|\leq\|\eta_{-}(u\oplus u^{\prime}_{\mathrm{path}})r-\eta_{-}(u^{\prime}_{\mathrm{amp}})r\|\leq\\ \leq C\|u\oplus u^{\prime}_{\mathrm{path}}-u^{\prime}_{\mathrm{amp}}\|\leq C\Big(7{\varepsilon}+C{\varepsilon}^{-1/2}\delta^{1/2}\Big)=:\gamma,$
where $C=\|\eta_{-}\|_{\text{OL}(\mathbb{T})}$ is an absolute constant.That is, the projections $1_{\Omega_{-}}(u^{\prime}_{\mathrm{amp}})$ and $r$ are approximately orthogonal, with the error coming from the transition among the unitaries $u\oplus u^{\prime}_{\mathrm{path}}$ and $u^{\prime}_{\mathrm{amp}}.$
Assuming that $\gamma<1/100,$ apply Corollary 2.9 to produce a unitary $\sigma\in A_{2d}$ such that
(4.5) $\sigma 1_{\Omega_{-}}(u^{\prime}_{\mathrm{amp}})\sigma^{*}\leq\mathbf{1}-r=p+q,\quad\left\|\sigma-\mathbf{1}\right\|\leq 5\gamma.$
So let
$w^{\dagger}:=\sigma u^{\prime}_{\mathrm{amp}}\sigma^{*},\quad v^{\dagger}:=\sigma v^{\prime}_{\mathrm{amp}}\sigma^{*},$
and notice that, up to constants, $w^{\dagger}$ and $v^{\dagger}$ satisfy the same assumptions as $u^{\prime}_{\mathrm{amp}}$ and $v^{\prime}_{\mathrm{amp}},$ but gain the property that the spectral subspace of $w^{\dagger}$, associated to $\Omega_{-}$, is contained in the range of $p+q$. We denote
$s_{-}:=1_{\Omega_{-}}(w^{\dagger})=\sigma 1_{\Omega_{-}}(u^{\prime}_{\mathrm{amp}})\sigma^{*}=(p+q)s_{-}.$
Since $p$ and $q$ commute with $u\oplus u^{\prime}_{\mathrm{path}}$ and therefore almost commutes with $u^{\prime}_{\mathrm{amp}}$ and $w^{\dagger}$, we have the following estimates for each of the commutators, either from equation (4.5) or from their construction:
(4.6) $\begin{rcases}\begin{aligned} \|[p,w^{\dagger}]\|,&\quad\|[q,w^{\dagger}]\|,\\ \|[p,v^{\dagger}]\|,&\quad\|[q,v^{\dagger}]\|,\\ \|[p,s_{-}]\|,&\quad\|[q,s_{-}]\|\end{aligned}\end{rcases}\leq C\gamma.$
Having made sufficient preparations, we now perform the first dimension reduction. In view of Proposition 2.11 and equation (4.6), we have that both $(p+q)w^{\dagger}|_{p+q}$ and $(p+q)v^{\dagger}|_{p+q}$ are $C\gamma$-close to some unitary operators on $\operatorname{ran}(p+q)$. Let us denote these operators by $g$ and $h$, and restrict our attention to $\operatorname{ran}(p+q)$, thus “forgetting” about $r$. We note that, since $s_{-}=(p+q)s_{-}$, it is preserved by the corresponding compression.
We now consider further partitioning $\operatorname{ran}(p+q)$ by $s_{-}$. By applying Proposition 2.11, again in view of equation (4.6), we have that each of the four operators
(4.7) $(p+q-s_{-})g|_{p+q-s_{-}},\quad s_{-}g|_{s_{-}},\quad(p+q-s_{-})h|_{p+q-s_{-}},\quad s_{-}h|_{s_{-}}$
is $C\gamma$-close to a unitary operator on the corresponding subspace. We denote these unitary operators by $g_{+},g_{-},h_{+},h_{-}$. From equation (4.6), they also almost commute up to $C\gamma$:
$\|[g_{+},h_{+}]\|+\|[g_{-},h_{-}]\|\leq C\gamma.$
Recall that, originally, $s_{-}$ was the spectral projection of $w^{\dagger}$ associated to $\Omega_{-}$, which is a fixed interval in $\mathbb{T}$. Therefore,
(4.8) $\sigma(s_{-}w^{\dagger}|_{s_{-}})\subset\Omega_{-},\quad\sigma((\mathbf{1}-s_{-})w^{\dagger}|_{\mathbf{1}-s_{-}})\subset\mathbb{T}\setminus\Omega_{-}.$
In other words, each of these operators has a large spectral gap, of diameter at least $1/5$. Claim (ii) of Proposition 2.11, applied twice for each of the restrictions (first for the compressions of $w^{\dagger}$ and $v^{\dagger}$, arriving to $g$ and $h$, and then for their compressions (4.7) arriving to to $_{+}, gg_{-}$, $_{+}, hh_{-}$), implies that $g_{+}$ and $g_{-}$ each also inherit spectral gaps of diameter, say, at least $1/5-C\gamma\geq 1/10$, assuming $\gamma<1/1000$.
We can now apply Lemma 4.2, with $\rho=1/10$, to the pairs $(g_{+},h_{+})$ and $(g_{-},h_{-})$. Thus, we derive two commuting pairs
$g_{+}^{\prime},h_{+}^{\prime}]=0,\quad[g_{-}^{\prime},h_{-}^{\prime}]=0,\quad\|g_{+}^{\prime}-g_{+}\|+\|g_{-}^{\prime}-g_{-}\|+\|h_{+}^{\prime}-h_{+}\|+\|h_{-}^{\prime}-h_{-}\|\leq C\gamma^{1/2}.$
Then, $g^{\prime}=g_{+}^{\prime}+g_{-}^{\prime}$ and $h^{\prime}=h_{+}^{\prime}+h_{-}^{\prime}$ are unitaries on $\operatorname{ran}(p+q)$, satisfying
(4.9) $g^{\prime},h^{\prime}]=0,\quad\|g^{\prime}-u\oplus w^{\prime}_{-}\|+\|h^{\prime}-v\oplus v^{\prime}_{-}\|\leq C\gamma^{1/2},\quad\text{where}\quad v^{\prime}_{-}:=qv^{\prime}_{\mathrm{path}}|_{q},\quad w^{\prime}_{-}=qu^{\prime}_{\mathrm{path}}|_{q}.$
4.5. Proof of Theorem 1.4: the second dimension reduction.
We now find ourselves in a situation somewhat similar to the previous Subsection 4.4. Similarly to disposing of $r$, our goal is now to dispose of $q$. Note that $\gamma$ is now replaced by $C\gamma^{1/2}$, but the rest is largely the same.
We start with commuting operators $g^{\prime}$ and $h^{\prime}$. By repeating the steps that led to equation (4.5), using the opposite semi-circle $\{\Re z\leq 0\}$ and the spectral projection $1_{\{\Re z>1/2\}}(g^{\prime})$, we generate another unitary $\tau\in U(\operatorname{ran}(p+q))$ with
(4.10) $\|\tau-\mathbf{1}_{\operatorname{ran}(p+q)}\|<100\gamma^{1/2}$
and
$s_{+}:=\tau 1_{\{\Re z>1/2\}}(g^{\prime})\tau^{*}\leq p,\quad g^{\dagger}:=\tau g^{\prime}\tau^{*},\quad h^{\dagger}:=\tau h^{\prime}\tau^{*}.$
From (4.9) and (4.10), we have
$\|g^{\dagger}-u\oplus w^{\prime}_{-}\|+\|h^{\dagger}-v\oplus v^{\prime}_{-}\|\leq C\gamma^{1/2}.$
Similarly to the previous subsection, the pair $(g^{\dagger},h^{\dagger})$ can now replace the pair $(g^{\prime},h^{\prime})$ for all practical purposes. With this replacement, we gained an additional property that $1_{\{\Re z>1/2\}}(g^{\dagger})\leq p$. We also have the estimates (4.6) with $w^{\dagger},v^{\dagger},s_{-},\gamma$ replaced by $g^{\dagger},h^{\dagger},s_{+},\gamma^{1/2}$, respectively. As a consequence, the compressions
$\quad(s_{+}g^{\dagger}|_{s_{+}},s_{+}h^{\dagger}|_{s_{+}})\quad\text{and}\quad((p-s_{+})g^{\dagger}|_{p-s_{+}},(p-s_{+})h^{\dagger}|_{p-s_{+}})$
form two pairs of almost unitary operators, with the first elements of each pair having large spectral gaps, similarly to equation (4.8), with the role of $\Omega_{-}$ now played by $\Omega_{+}=\{\Re z>1/2\}\cap\mathbb{T}$.
More precisely, we can apply Proposition 2.11 with $\varkappa=\rho=C\gamma^{1/2}\ll 1$ and, in the case of the compressions of $g^{\dagger}$, with the gap diameter $R=1/20$ and center $z=1$ for the compression associated to $s_{+}$ and z=1z=-1 for that corresponding to $(p-s_{+})$.
As a result, we obtain two pairs of exactly unitary operators, but lost exact commutation. Each pair is now $C\gamma^{1/2}$-almost commuting, and the first element of each pair has a gap of diameter, say, $1/30$. This allows another, final, application of Lemma 4.2 to each pair, resulting in two commuting pairs of unitaries $(u^{\prime}_{+},v^{\prime}_{+})$ on $\operatorname{ran}(s_{+})$ and $(u_{-}^{\prime},v^{\prime}_{-})$ on $\operatorname{ran}(p-s_{+})$, which sum to the final commuting pair
$u^{\prime}:=u^{\prime}_{+}+u^{\prime}_{-},\quad v^{\prime}:=v^{\prime}_{+}+v^{\prime}_{-}$
with
$\|u^{\prime}-u\|+\|v^{\prime}-v\|<C\gamma^{1/4}.$
The commuting unitary matrices $u^{\prime}$, $v^{\prime}$, together with the estimate above, conclude the proof of Theorem 1.4 by noting that
$\varepsilon=\delta^{1/3},\quad\gamma=C\delta^{1/3}$
provides the balanced choice of the parameters leading to the distance estimate $C\|[u,v]\|^{1/12}$.∎
Appendix A Operator Lipschitz functions
In this section, we will discuss the definition and some properties of operator Lipschitz functions. Let $\mathcal{F}\subset\mathbb{C}$ be a closed subset, and $f\colon\mathcal{F}\to\mathbb{C}$ be a continuous function. We say that $f\in\operatorname{OL}(\mathcal{F})$ (that is, operator Lipschitz on $\mathcal{F}$) if, for some $C>0$, we have
(A.1) $\|f(A_{1})-f(A_{2})\|\leq C\|A_{1}-A_{2}\|$
for all bounded normal normal operators $A_{1},A_{2}\in B(H)$ with $\sigma(A_{j})\in\mathcal{F}$. The smallest possible value of $C$ in equation (A.1) will be denoted by $\|f\|_{\operatorname{OL}(\mathcal{F})}$. It is easy to see that $\|\cdot\|_{\operatorname{OL}(\mathcal{F})}$ is a seminorm on $\operatorname{OL}(\mathcal{F})$, that vanishes only on constant functions. Most commonly, one considers $\mathcal{F}=\mathbb{C},\mathbb{R},\mathbb{T}$, which corresponds to functions defined on normals, self-adjoint, and unitary operators, respectively. We refer the reader to [1, 2, 3] for a comprehensive review of the properties of operator Lipschitz functions, with proofs and references. In what follows, we provide a summary of the properties that are used in the present paper. We start from some basic facts.
  • (OL1)
    Every operator Lipschitz function is Lipschitz, with Lipschitz constant equal to $\|f\|_{\operatorname{OL}(\mathcal{F})}$. The converse is not necessarily true: for example, $|\cdot|\notin\|f\|_{\operatorname{OL}(\mathbb{R})}$.
  • (OL2)
    $\operatorname{OL}(\mathcal{F})/\mathbb{C}$ (that is, operator Lipschitz functions modulo constant functions) is a Banach space.
  • (OL3)
    Let $C_{b}^{2}(\mathbb{C})$ be the space of bounded continuously twice differentiable functions with bounded first and second partial derivatives. Then $C_{b}^{2}(\mathbb{C})\subset\operatorname{OL}(\mathbb{C})$, and $\|f\|_{\operatorname{OL}(\mathbb{C})}\leq C\|f\|_{C_{b}^{2}(\mathbb{C})}$. Note that the same also holds for $\mathbb{C}$ replaced by $\mathcal{F}$, if one defines $C_{b}^{2}(\mathcal{F})$ to be the set of functions that admit an extension in $C_{b}^{2}(\mathbb{C})$.
  • (OL4)
    A linear function $f(z)=az+b$ is operator Lipschitz on $\mathbb{C}$, even though it does not belong to $C_{b}^{2}(\mathbb{C})$.
  • (OL5)
    Let $a,b\in\mathbb{C}$, $a\neq 0$. Define an affine transformation $t\colon\mathbb{C}\to\mathbb{C}$ by $t(z):=az+b$. The operator Lipschitz norm behaves in the same way (and with the same proof) as the Lipschitz norm under such transformations. That is,
    $\|f\circ t\|_{\operatorname{OL}(t^{-1}(\mathcal{F}))}=|a|\|f\|_{\operatorname{OL}(\mathcal{F})}.$
  • (OL6)
    Every operator Lipschitz function is commutator Lipschitz. That is,
    $\|[f(N),B]\|\leq\|f\|_{\operatorname{OL}(\mathcal{F})}\|[N,B]\|$
    for every normal operator $N$ with $\sigma(N)\subset\mathcal{F}$ and every $B$ that is a unitary or a bounded self-adjoint.
  • (OL7)
    Let $f\colon\mathbb{R}\to\mathbb{C}$ be $1$-periodic. Then $f\in\operatorname{OL}(\mathbb{R})$ if and only if $f(t)=g(e^{2\pi it})$, for some $g\in\operatorname{OL}(\mathbb{T})$. In other words, a $1$-periodic function is operator Lipschitz on $\mathbb{R}$ if and only if the associate function on the circle is operator Lipschitz on $\mathbb{T}$.
  • (OL8)
    Suppose that $f,g\in\operatorname{OL}(\mathcal{F})\cap L^{\infty}(\mathcal{F})$ are bounded operator Lipschitz functions. Then $fg\in\operatorname{OL}(\mathcal{F})$ and
    $\|fg\|_{\operatorname{OL}(\mathcal{F})}\leq\|f\|_{L^{\infty}(\mathcal{F})}\|g\|_{\operatorname{OL}(\mathcal{F})}+\|f\|_{\operatorname{OL}(\mathcal{F})}\|g\|_{L^{\infty}(\mathcal{F})}.$
Let $\mathbb{T}=\{z\in\mathbb{C}\colon|z|=1\}$ be the unit circle, and
$\mathbb{T}_{\rho}:=\{z\in\mathbb{C}\colon|z+1|\geq\rho\},\quad\rho>0,$
be the same circle with a small arc of distance $\rho$ from 1-1 removed. Denote by $\arg_{-}$ the branch of the complex argument function that maps $\mathbb{C}\setminus(-\infty,0]$ continuously onto $(-\pi,\pi)$. The following estimate will be important in order to avoid additional commutator norm losses in the proof of the main result.
Lemma A.1.
There is an absolute constant $C>0$ such that, for $0<\rho<2$, we have
$\|\arg_{-}\|_{\operatorname{OL}(\mathbb{T}_{\rho})}\leq C\rho^{-1}.$
Proof.
Clearly, (OL3) implies a similar statement with $C\rho^{-2}$ in the right hand side. As a consequence, it is sufficient to prove the lemma for, say, $0<\rho<1/100$. Let $\arg_{+}\colon\mathbb{C}\setminus[0,+\infty)\to(0,2\pi)$ be another branch of the complex argument that is smooth near 1-1. It is easy to see that
(A.2) $\arg_{-}(z)=\arg_{+}(z)-\pi(1-\operatorname{sign}y),\quad z=x+iy,\quad|z+1|<1.$
We will construct a smooth function on $\mathbb{C}$ that coincides with $\arg_{-}$ on $\mathbb{T}_{\rho}$ and whose operator Lipschitz norm we can control. In order to do so, let $s\colon\mathbb{R}\to[-1,1]$ be a smooth function such that
$s(x)=\begin{cases}-1,&x\leq-1;\\ 1,&x\geq 1.\end{cases}$
Let also
(A.3) $s_{\rho}(z)=s_{\rho}(x+iy):=s(2y/\rho).$
From (OL3) and (OL5), it follows that
$\|s_{\rho}\|_{\operatorname{OL}(\mathbb{C})}\leq C\rho^{-1},$
where $C$ is an absolute constant. Let $\varphi,\psi\colon\mathbb{C}\to[0,1]$ be smooth compactly supported functions such that
$\psi(z)=\begin{cases}1,&z\in\mathbb{T};\\ 0,&\operatorname{dist}(z,\mathbb{T})\geq 1/10;\end{cases}\quad\varphi(z)=\begin{cases}1,&|z+1|\leq 1/10;\\ 0,&|z+1|\geq 1/5.\end{cases}$
Finally, define
(A.4) $\arg_{\rho}(z):=\varphi(z)(\arg_{+}(z)+\pi(1-s_{\rho}(z)))+(1-\varphi(z))\psi(z)\arg_{-}(z).$
The reader can easily check that $\arg_{\rho}(z)=\arg_{-}(z)$ for $z\in\mathbb{T}_{\rho}$, since
$s_{\rho}(z)=s_{\rho}(x+iy)=\operatorname{sign}(y),\quad z\in\mathbb{T}_{\rho}\cap\mathrm{supp}\,\varphi.$
We have
$\|\arg_{\rho}\|_{\operatorname{OL}(\mathbb{C})}\leq\|\varphi\arg_{+}-\pi\varphi\|_{\operatorname{OL}(\mathbb{C})}+\|(1-\varphi)\psi\arg_{-}\|_{\operatorname{OL}(\mathbb{C})}+\pi\|\varphi s_{z}\|_{\operatorname{OL}(\mathbb{C})}.$
Since $\arg_{+}$ is smooth on $\mathrm{supp}\,\varphi$ and $\arg_{-}$ is smooth on $\mathrm{supp}\,(1-\varphi)\psi$, Property (OL3) implies that the first two terms are bounded by absolute constants. The third term is bounded by $C^{\prime}\rho^{-1}$ due to (A.3), (OL8), and (OL3), where $C^{\prime}$ is another absolute constant. ∎